CS 2001: Research Topics in Computer Science

Fall 2013

Mini Project

Assigned: Thursday, October 10 Due: Thursday, December 12

1 Background

The goal of this short project is to give you a working knowledge of AFS, Subversion (SVN), make, and L^{A}TEX. Specifically, you will learn how to create groups and manage permissions in AFS, set up an SVN repository whose permissions are managed via AFS groups, typeset documents using multiple L^{A}TEX styles, and control document builds using Makefiles.

2 What do I need to do?

To address the above objectives, you will set up an SVN repository to share your paper reviews with me. Given that everyone in the department has access to the shared AFS file system, this process is actually quite easy to do in practice. In particular, you will need to do the following things:

1. Create an AFS group called cs2001 svn. After the group is created, add yourself and me (znati) as members of the group.

2. Within your public AFS space, set up an SVN repository called cs2001 papers. Ensure that only members of the cs2001 svn group have access to this repository.

3. Read over the paper submission instructions for the 2013 ACM Conference on Computer and Communications Security (http://www.sigsac.org/ccs/CCS2013/) and the 2014 IEEE Symposium on Security and Privacy (http://www.ieee-security. org/TC/SP2014/). Each of these conferences requires the use of a different L^{A}TEX template; download both templates.

4. Create two directories in your new SVN repository: 15 nov and 22 nov. In the 15 nov directory, create a L^{A}TEX document that formats your paper review that was due on November 15th according to the 2013 ACM Conference on Computer and Communications Security format. In the 22 nov directory, create a L^{A}TEX document that formats your paper review that was due on November 22nd according to the 2014 IEEE Symposium on Security and Privacy format. Each review should have three sections: Summary, Technical Strengths, and Technical Weaknesses.

5. In each review, the paper that you chose to review should be cited. Do not create the L^{A}TEX bibliography by hand. Rather, create a BibTeX file review.bib that includes the citation information for this paper. Your L^{A}TEX document should use the \cite{} command to include a citation to the paper defined in review.bib.

6. In each directory, create a Makefile with two targets: review and clean. The review target should use pdflatex and bibtex to build the paper review in the current directory. The resulting PDF file should include a references section that cites (at least) the paper that you reviewed. The clean target should delete all auxiliary files created by the pdflatex and bibtex programs. That is, after running make clean, the only things in the directory should be Makefile, your .tex, .bib, and .cls files.

The resulting SVN repository will allow me to download your reviews, edit them to include comments and suggestions, and commit my changes to the repository. In order to carry out the above, you may need to consult references on the course web page, the department’s web page, or other places on the Internet.

3 What should I turn in?

When you have the above SVN repository set up, send me (znati@cs.pitt.edu) an email containing:

• Instructions on how to check out your SVN repository. I should be able to copy and paste these instructions directly into a command prompt to check out a copy of your

cs2001 papers repository.

• The membership list of your cs2001 svn AFS group. Specifically, I expect to see a copy/paste showing the results of an AFS command that lists group membership.

• The AFS permissions for your repository. Again, I expect to see a copy/paste showing the results of an AFS command that lists the permissions on the cs2001 papers directory.

Please email all of the above information to znati@cs.pitt.edu by noon on Thursday December 12th, 2013.